OpenStudy (meilendurcer):

Suppose you rent a limousine for a formal reception. The bill for the evening is $53.00. A tax of 6% will be added, and you want to tip the chauffeur 20% for excellent driving. How much will you pay in total?

OpenStudy (anonymous):

53* .06=3.18 and 53*.2=10.6 So you would pay $66.78. I don't believe you would factor in the tax to the drivers tip however if you do then you would simply do 56.18*.2.

OpenStudy (anonymous):

About 67 bucks. ($66.78 rounded up, boy, am I a sport!)

OpenStudy (meilendurcer):

Thank you!!

OpenStudy (anonymous):

The calculation was 53 * 1.26.

OpenStudy (meilendurcer):

ahh ok..

OpenStudy (lilg132):

53/100 = 0.53 (THIS IS 1% OF THE PRICE) WE WANT TO WORK OUT HOW MUCH 6% WILL BE OF THIS SO WE MULTIPLY BY 6 0.53 * 6 = 3.18 WE ADD 3.18 TO THE PRICE 53 + 3.18 = 56.18 NOW TO GIVE HIM 20% TIP OF THAT PRICE WE DO THE SAME THING 56.18 / 100 = 0.5618 (THIS IS 1% OF THE PRICE) TO WORK OUT 20% WE MULTIPLY BY 20 0.5618 * 20 = 11.2316 NOW WE ADD THIS TO THE TOTAL PRICE 56.18 + 11.2316 = 67.416 AND BECAUSE WE ARE WORKING WITH A CURRENCY WE HAVE TO USE 2 DECIMAL PLACES SO THE FINAL BILL WILL BE $67.42 DONT FORGET TO CLICK BEST RESPONSE

OpenStudy (lilg132):

also it doesnt say to round up so your answer would be wrong if you rounded it up to $67

OpenStudy (meilendurcer):

right and sometimes they mix it up to make it different

OpenStudy (lilg132):

yes please dont forget to click best response :)

OpenStudy (anonymous):

Tips are usually calculated on the charges, not on the combined charges and tax. I think the exact answer the author of your text expected is the $66.78.

OpenStudy (lilg132):

i come from england where tax is always included with the price my bad :p

OpenStudy (anonymous):

No sweat. Do math every day.